# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

### v2

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message passing) layers. Within each of these layers, the representation of each node is updated from an aggregation and transformation of its neighbours representations at the previous layer. The upper bound for the expressive power of message passing GNNs was reached through the use of MLPs as a transformation, due to their universal approximation capabilities. However, MLPs suffer from well-known limitations, which recently motivated the introduction of Kolmogorov-Arnold Networks (KANs). KANs rely on the Kolmogorov-Arnold representation theorem, rendering them a promising alternative to MLPs. In this work, we compare the performance of KANs against that of MLPs in graph learning tasks. We perform extensive experiments on node classification, graph classification and graph regression datasets. Our preliminary results indicate that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks. Code is available at https://github.com/RomanBresson/KAGNN.
@@ 1 Introduction
Graphs are structural representations of information which are useful for modeling many types of data. They arise naturally in a wide range of application domains, and their abstract nature offers increased flexibility. Typically, the nodes of a graph represent entities, while the edges capture the interactions between them. For instance, in social networks, nodes represent individuals, and edges represent their social interactions. In chemo-informatics, molecules are commonly modeled as graphs, with nodes corresponding to atoms and edges to chemical bonds. In other settings, molecules can also be nodes, with edges capturing their ability to bond with one another.
@@
In many cases where graph data is available, there exist problems that cannot be solved efficiently using conventional tools (e. g., graph algorithms) and require the use of machine learning techniques. For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations [12]. Machine learning methods could serve as a more efficient alternative to those methods. Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs [36]. Most GNNs consist of a series of message passing layers. Within a message passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
@@
A lot of recent work has focused on investigating the expressive power of GNNs [22]. There exist different definitions of expressive power, however, the most common definition is concerned with the number of pairs of non-isomorphic graphs that a GNN model can distinguish. Two graphs are isomorphic if there exists an edge-preserving bijection between their respective sets of nodes. In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot [27]. Furthermore, an equivalence has also been established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [3]. This line of work gave insights into the limitations of different models [38, 29], but also led to the development of more powerful architectures [27, 24, 26].
@@
Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities [5, 15]. The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons. Having said that, in practice the models suffer from several limitations due to non-convex loss functions, algorithms without convergence guarantees and a notorious lack of interpretability that hinders their applicability in several domains. Recently, Kolmogorov-Arnold Networks (KANs) [23] have emerged as promising alternatives to MLPs. They are based on the Kolmogorov-Arnold representation theorem [21] which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions. KANs substitute the learnable weights and pre-defined activation functions of MLPs, with learnable activations based on B-splines and summations. The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions, while simultaneously being more interpretable.
@@
In this paper, we present a thorough empirical comparison between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end. Our work is orthogonal to prior work that studies the expressive power of GNNs. Here, we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i. e., KANs or MLPs) on the model’s performance. We evaluate the different GNN models on several standard node classification, graph classification and graph regression datasets.
@@
The rest of this paper is organized as follows. Section 2, provides an overview of the tasks we address in this paper, as well as a description of message passing GNNs and Kolmogorov-Arnold networks. In Section 3, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network) and KAGCN (Kolmogorov-Arnold Graph Convolution Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer. In Section 4, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks. Finally, Section 5 concludes the paper.
@@ 2 Background
2.1 Considered Graph Learning Tasks
Before presenting the tasks on which we focus in this study, we start by introducing some key notation for graphs. Let $\mathbb{N}$ denote the set of natural numbers, i. e., $\{1,2,\ldots\}$. Then, $[n]=\{1,\ldots,n\}\subset\mathbb{N}$ for $n\geq 1$. Let $G=(V,E)$ be an undirected graph, where $V$ is the vertex set and $E$ is the edge set. We denote by $n$ the number of vertices and by $m$ the number of edges, i. e., $n=|V|$ and $m=|E|$. Let $g\colon V\rightarrow[n]$ denote a bijective mapping from the space of nodes to set $[n]$. Let $\mathcal{N}(v)$ denote the the neighbourhood of vertex $v$, i. e., the set $\{u\mid\{v,u\}\in E\}$. The degree of a vertex $v$ is $\deg(v)=|\mathcal{N}(v)|$. Each node $v\in V$ is associated with a $d$-dimensional feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$. Thus, $\mathbf{x}_{v}$ is equal to the $g(v)$-th row of $\mathbf{X}$.
In node classification, each node $v\in V$ is associated with a label $y_{v}$ that represents a class. The task is to learn a function that maps nodes to their class labels, i. e., to learn a function $f_{\text{node}}$ such that $f_{\text{node}}(v,G,\mathbf{X})=y_{v}$. In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$. The task is then to learn a function that maps graphs to their class labels/targets, i. e., to learn a function $f_{\text{graph}}$ such that $f_{\text{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space. That is, we aim to learn $\mathbf{H}=\texttt{EMBEDDING}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension. Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$. Let $\mathbf{h}_{v}$ denote this embedding. For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$. For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=\texttt{READOUT}(\mathbf{H})$ for the entire graph.
One particularly desirable property of such models is permutation invariance. That is, the embedding $\mathbf{h}_{G}$ of a graph needs to be the same regardless of the ordering of its nodes . Indeed, these orderings do not hold any semantic meaning and different orderings give rise to isomorphic graphs. Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max or mean operators.
@@ 2.2 Graph Neural Networks
One of the most widely-used paradigms for designing such permutation invariant models is the message passing framework [12] which consists of a sequence of layers and whithin each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings. Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
$\mathbf{h}_{v}^{(\ell)}=\phi^{(l)}\left(\mathbf{h}_{v}^{(\ell-1)},\bigoplus\limits_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(\ell-1)}\right)$ (1)
@@
As discussed above, in this paper, we focus on the functions that different GNN models employ to update node representations. Many existing GNNs use a $1$-layer perceptron (i. e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features [7, 20, 40]. For instance, each layer of the the Graph Convolutional Network (GCN) [20] is defined as follows:
@@
However, the 1-layer perceptron is not a universal approximator of multiset functions [38]. Thus, the emerging GNN might not be expressive enough for some tasks. Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations [2, 6, 28, 29]. It is well-known that standard message passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [38]. While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
@@
The GIN model can achieve its full potential if proper weights (i. e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned. However, in practice, GIN might fail to learn those weights due to limited training data and due to limitations of the employed training algorithm (e. g., stochastic gradient descent). This has motivated a series of works which focused on improving the training procedure of GNNs. For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training [13]. Other works have studied how to initialize the weights of the MLPs of the message passing layers of GNNs. It was shown that by adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [14]. On the other hand, there exist settings where there is no need of complex learning models. This has led to the development of methods for simplifying GNNs. This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [35] or by feeding the node features into a neural network which generates predictions and then propagate those predictions via a personalized PageRank scheme [9].
@@ 2.3 Kolmogorov-Arnold Networks
where all $\Phi_{\square}$ and $\phi_{\square}$ functions are univariate, and the sum is the only multivariate operator.
Equation (4) can be seen as a two-step process. First, a different set of univariate non-linear activation functions is applied to each dimension of the input, and then the output of those functions are summed up. The authors rely on this interpretation to define a Kolmogorov-Arnold Network (KAN) layer, which is a mapping between a space $A\subseteq\mathbb{R}^{d}$ and a different space $B\subseteq\mathbb{R}^{d^{\prime}}$, (identical in use to an MLP layer). Such a layer consists of $d\times d^{\prime}$ trainable functions $\{\phi_{ij},~{}1\leq i\leq d^{\prime},~{}1\leq j\leq d\}$. Then, for $\mathbf{x}\in A$, we compute its image $\mathbf{x}^{\prime}$ as:
@@
Stacking two such layers, one with input dimension $d$ and output dimension $12d+$ 1, and another with input dimension $12d+$ 1 and output dimension $1$, we obtain Equation (4), and the derived model is a universal function approximator.
This seemingly offers a complexity advantage compared to MLPs, since the number of univariate functions required to represent any multivariate function from $\left[0,1\right]^{d}$ to $\mathbb{R}^{d^{\prime}}$ is at most $(2d^{2}+d)\times d^{\prime}$, whereas the universal approximation theorem for the MLP requires a possibly infinite number of neurons. However, as stated by the original paper, the behavior of such univariate functions might be arbitrarily complex (e. g., fractal, non-smooth), thus leading to them being non-representable, and non-learnable.
MLPs relax the arbitrary-width constraint by stacking finite-width layers. Likewise, KANs relax the arbitrary-complexity constraints on the non-linearities by stacking KAN layers. Thus, the output of a function is given by:
$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})$ (6)
where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.
The original paper uses splines (i. e., trainable piecewise-polynomial functions) as nonlinearities. This allows to retain a high expressivity for a relatively small number of parameters, at the cost of enforcing some local smoothness. A layer $\ell$ is thus a $d_{\ell}\times d_{\ell-1}$ grid of splines. The degree used for each spline (called spline order), as well as the number of splines used for each function (called grid size) are both hyperparameters of the architecture.
Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification [4] and for predicting the pressure and flow rate of flexible electrohydrodynamic pumps [31]. So far, most efforts have focused on time series data [39]. For instance, KANs have been evaluated in the satellite traffic forecasting task [34]. Furthermore, they have been combined with architectures that are traditionally leveraged in time series forecasting tasks such as the Long Short-Term Memory Network [11] and the Transformer [10]. The work closest to ours is the one reported in [37], where the authors propose FourierKAN-GCF. This is a GNN model designed for the task of graph collaborative filtering where the feature transformation in the neighborhood aggregation layers is performed by KANs.

### v3

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message-passing) layers, within which the representation of each node is updated based on those of its neighbors. The most expressive message-passing GNNs can be obtained through the use of the sum aggregator and of MLPs for feature transformation, thanks to their universal approximation capabilities. However, the limitations of MLPs recently motivated the introduction of another family of universal approximators, called Kolmogorov-Arnold Networks (KANs) which rely on a different representation theorem. In this work, we compare the performance of KANs against that of MLPs on graph learning tasks. We evaluate two different implementations of KANs using two distinct base families of functions, namely B-splines and radial basis functions. We perform extensive experiments on node classification, graph classification and graph regression datasets. Our results indicate that KANs are on-par with or better than MLPs on all studied tasks, making them viable alternatives, at the cost of some computational complexity. All code will be made available upon acceptance.
@@ 1 Introduction
Graphs are structural representations of information useful for modeling different types of data. They arise naturally in a wide range of application domains, and their abstract nature offers increased flexibility. Typically, the nodes of a graph represent entities, while the edges capture the interactions between them. For instance, in social networks, nodes represent individuals, and edges represent their social interactions. In chemo-informatics, molecules are commonly modeled as graphs, with nodes corresponding to atoms and edges to chemical bonds. In other settings, molecules can also correspond to nodes, with edges capturing their ability to bond with one another.
@@
In many cases where graph data is available, there exist problems that cannot be solved efficiently using conventional tools (e. g., graph algorithms) and require the use of machine learning techniques. For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations [1]. Machine learning methods could serve as a more efficient alternative to those methods. Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs [2]. Most GNNs consist of a series of message-passing layers. Within a message-passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
@@
A lot of recent work has focused on investigating the expressive power of GNNs [3]. There exist different definitions of expressive power, however, the most common definition is concerned with the number of pairs of non-isomorphic graphs that a GNN model can distinguish. Two graphs are isomorphic if there exists an edge-preserving bijection between their respective sets of nodes. In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot [4]. Furthermore, an equivalence was established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [5]. This line of work gave insights into the limitations of different models [6, 7], but also led to the development of more powerful architectures [4, 8, 9].
@@
Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities [10, 11]. The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons. Recently, Kolmogorov-Arnold Networks (KANs) [12] have emerged as promising alternatives to MLPs. They are based on the Kolmogorov-Arnold representation theorem [13] which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions. KANs substitute the learnable weights and pre-defined activation functions of MLPs with learnable activations and summations. The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions and in settings where regularity is expected.
@@
In this paper, we present a thorough empirical comparison of the performance between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end. Our intuition is that KANs could exploit the expected regularity underlying the message-passing paradigm to provide comparatively good performance on graph tasks. Our work is orthogonal to prior work that studies the expressive power of GNNs since we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i. e., KANs or MLPs) on the models’ performance. We evaluate the different GNN models on several standard node classification, graph classification, and graph regression datasets.
@@
The rest of this paper is organized as follows. Section 2, provides an overview of the tasks we address in this paper, as well as a description of message-passing GNNs and Kolmogorov-Arnold networks. In Section 3, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network) and KAGCN (Kolmogorov-Arnold Graph Convolution Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer. In Section 4, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks. Finally, Section 7 concludes the paper.
@@ 2 Background
2.1 Graph Learning Tasks
Before presenting the tasks on which we focus in this study, we start by introducing some key notation for graphs. Let $\mathbb{N}$ denote the set of natural numbers, i. e., $\{1,2,\ldots\}$. Then, $[n]=\{1,\ldots,n\}\subset\mathbb{N}$ for $n\geq 1$. Let $G=(V,E)$ be an undirected graph, where $V$ is the vertex set and $E$ is the edge set. We denote by $n$ the number of vertices and by $m$ the number of edges, i. e., $n=|V|$ and $m=|E|$. Let $g\colon V\rightarrow[n]$ denote a bijective mapping from the space of nodes to set $[n]$. Let $\mathcal{N}(v)$ denote the neighbourhood of vertex $v$, i. e., the set $\{u\mid\{v,u\}\in E\}$. The degree of a vertex $v$ is $\deg(v)=|\mathcal{N}(v)|$. Each node $v\in V$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$. Thus, $\mathbf{x}_{v}$ is equal to the $g(v)$-th row of $\mathbf{X}$.
In node classification, each node $v\in V$ is associated with a label $y_{v}$ that represents a class. The task is to learn how to map nodes to their class labels, i. e., to learn a function $f_{\texttt{node}}$ such that $f_{\texttt{node}}(v,G,\mathbf{X})=y_{v}$. In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$. The task is then to learn a function that maps graphs to their class labels/targets, i. e., a function $f_{\texttt{graph}}$ such that $f_{\texttt{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space. That is, we aim to learn $\mathbf{H}=f_{\texttt{embedding}}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension. Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$. Let $\mathbf{h}_{v}$ denote this embedding. For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$. For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=f_{\texttt{readout}}(\mathbf{H})$ for the entire graph. One particularly desirable property of such models is permutation invariance. That is, the embedding $\mathbf{h}_{G}$ of a graph needs to be the same regardless of the ordering of its nodes. Indeed, these orderings do not hold any semantic meaning, and different orderings give rise to isomorphic graphs. Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max, or mean operators.
@@ 2.2 Graph Neural Networks
One of the most widely-used paradigms for designing such permutation invariant models is the message-passing framework [1] which consists of a sequence of layers and within each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings. Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
$\mathbf{h}_{v}^{(\ell)}=\phi^{(\ell)}\left(\mathbf{h}_{v}^{(\ell-1)},\bigoplus\limits_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(\ell-1)}\right)$ (1)
@@
As discussed above, we focus here on the functions that different GNN models employ to update node representations. Some GNNs use a $1$-layer perceptron (i. e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features [14, 15, 16]. For instance, each layer of the Graph Convolutional Network (GCN) [15] is defined as follows:
@@
However, the 1-layer perceptron is not a universal approximator of multiset functions [6], limiting the expressivity of the model. Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations [17, 18, 19, 7]. It is well-known that standard message-passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [6]. While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
@@
The GIN model can achieve its full potential if proper weights (i. e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned, which is not guaranteed. This has motivated a series of works that focused on improving the training procedure of GNNs. For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training [20]. Other works have studied how to initialize the weights of the MLPs of the message-passing layers of GNNs. It was shown that adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [21]. On the other hand, there exist settings where there is no need for complex learning models. This has led to the development of methods for simplifying GNNs. This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [22] or by feeding the node features into a neural network which generates predictions and then propagating those predictions via a personalized PageRank scheme [23].
@@ 2.3.1 Principle
where all $\Phi_{\square}$ and $\phi_{\square}$ denote univariate functions, and the sum is the only multivariate operator.
Equation 4 can be seen as a two-step process. First, a different set of univariate non-linear activation functions is applied to each dimension of the input, and then the output of those functions are summed up. The authors rely on this interpretation to define a Kolmogorov-Arnold Network (KAN) layer, which is a mapping between a space $A\subseteq\mathbb{R}^{d}$ and a different space $B\subseteq\mathbb{R}^{d^{\prime}}$, (identical in use to an MLP layer). Such a layer consists of $d\times d^{\prime}$ trainable functions $\{\phi_{ij},~{}1\leq i\leq d^{\prime},~{}1\leq j\leq d\}$. Then, for $\mathbf{x}\in A$, we compute its image $\mathbf{x}^{\prime}$ as:
@@
Stacking two such layers, one with input dimension $d$ and output dimension $12d+$ 1, and another with input dimension $12d+$ 1 and output dimension $1$, we obtain Equation 4, and the derived model is a universal function approximator. This seemingly offers a complexity advantage compared to MLPs, since the number of univariate functions required to represent any multivariate function from $\left[0,1\right]^{d}$ to $\mathbb{R}^{d^{\prime}}$ is at most $(2d^{2}+d)\times d^{\prime}$, whereas the universal approximation theorem for the MLP requires a possibly infinite number of neurons. However, as stated in the original paper, the behavior of such univariate functions might be arbitrarily complex (e. g., fractal, non-smooth), thus leading to them being non-representable and non-learnable.
MLPs relax the infinite-width constraint by stacking finite-width layers. Likewise, KANs relax the arbitrary complexity constraints on the non-linearities by stacking KAN layers of lower complexity. Thus, the output of a function is given by:
$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})~{}~{}\texttt{ where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.}$. (6)
2.3.2 KAN Implementations and Applications
The original paper that introduced KANs uses splines (i. e., trainable piecewise-polynomial functions) as nonlinearities. This allows to retain a high expressivity for a relatively small number of parameters, at the cost of enforcing some local smoothness. A layer $\ell$ is thus a $d_{\ell}\times d_{\ell-1}$ grid of splines. The degree used for each spline (called spline order), as well as the number of splines used for each function (called grid size) are both hyperparameters of the architecture. Nonetheless, KANs allow for any function to serve as its base element. Thus, alternative implementations use different bases, such as radial-basis functions [24], which in particular alleviate some computational bottleneck of splines.
Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification [25] and for predicting the pressure and flow rate of flexible electrohydrodynamic pumps [26]. So far, most efforts have focused on time series data [27]. For instance, KANs have been evaluated in the satellite traffic forecasting task [28]. Furthermore, they were combined with architectures usually leveraged in time series forecasting tasks such as the LSTM Network [29] and the Transformer [30]. The work closest to ours is the one reported in [31], where the authors propose FourierKAN-GCF. This is a GNN model designed for the task of graph collaborative filtering where the feature transformation in the neighborhood aggregation layers is performed by KANs.